# The Translation Trap Hypothesis for Multilingual Generation with Large Language Models The Translation Trap Hypothesis: Translation as a Constraint on Multilingual Generation

### v1

@@ Acknowledgments
We would like to thank Kaiser Sun and Krithika Ramesh for their helpful feedback. This work is partially supported by the AI and Advanced Computing Institute at Schmidt Sciences.
@@ Appendix A Languages
| Language | Language (Sub)Family | FLORES+ code | Supported by Aya | Supported by Llama |
| English | Germanic | eng_Latn | ✓ | ✓ |
| Spanish | Romance | spa_Latn | ✓ | ✓ |
| French | Romance | fra_Latn | ✓ | ✓ |
| German | Germanic | deu_Latn | ✓ | ✓ |
| Russian | East Slavic | rus_Cyrl | ✓ | - |
| Japanese | Japonic | jpn_Jpan | ✓ | - |
| Portuguese | Romance | por_Latn | ✓ | ✓ |
| Italian | Romance | ita_Latn | ✓ | ✓ |
| Turkish | Turkic | tur_Latn | ✓ | - |
| Korean | Koreanic | kor_Hang | ✓ | - |
| Arabic | Semitic | arb_Arab | ✓ | - |
| Polish | West Slavic | pol_Latn | ✓ | - |
| Hindi | Indic | hin_Deva | ✓ | ✓ |
| Hebrew | Semitic | heb_Hebr | ✓ | - |
| Ukranian | East Slavic | ukr_Cyrl | ✓ | - |
| Chinese (Simplified) | Sinitic | zho_Hans | ✓ | - |
| Chinese (Traditional) | Sinitic | zho_Hant | ✓ | - |
| Vietnamese | Austroasiatic | vie_Latn | ✓ | - |
| Indonesian | Austronesian | ind_Latn | ✓ | - |
| Romanian | Romance | ron_Latn | ✓ | - |
| Dutch | Germanic | nld_Latn | ✓ | - |
| Czech | West Slavic | ces_Latn | ✓ | - |
| Greek | Hellenic | ell_Grek | ✓ | - |
| Marathi | Indic | mar_Deva | - | - |
| Swahili | Niger-Congo | swh_Latn | - | - |
| Nepali | Indo-Aryan | nep_Deva | - | - |
| Tamil | Southern Dravidian | tam_Taml | - | - |
| Telugu | Southern Dravidian | tel_Telu | - | - |
| Uzbek | Turkic | uzn_Latn | - | - |
| Catalan | Romance | cat_Latn | - | - |
| Bosnian | South Slavic | bos_Latn | - | - |
| Cebuano | Austronesian | ceb_Latn | - | - |
| Persian | Indo-Iranian | pes_Arab | ✓ | - |
| Thai | Kra-Dai | tha_Thai | - | ✓ |
| Amharic | Semitic | amh_Ethi | - | - |
| Yoruba | Niger-Congo | yor_Latn | - | - |
@@ Appendix B Dataset Details
In order to curate the source concepts that we translate to every language, we prompted ChatGPT to give us instances each of English nouns, verbs, adjectives and adverbs. We then filtered out any words that did not appear in the English NLTK WordNet Bird and Loper (2004); Miller (1994), and manually vetted the remaining words to retain “general” concepts (i.e. excluding any proper nouns as well as highly specific cultural terms). We also checked for offensive words and did not encounter any. Finally, we sample words per part-of-speech of the filtered words, and conduct all our experiments on the resulting concepts.
We collect translations, including synonyms and alternate word senses, for these concepts for each of the other languages studied using Google Translate as mentioned in § 2. For average number of equivalent words per source word per language, see Table 3.
@@
The Turkish annotation was performed by an author on the paper, and the Polish annotation by a volunteer (acknowledged in the final version). Both annotators had native or near-native fluency. The consent of the annotators was obtained regarding the usage of the collected data for accuracy and coverage reporting of the lexica. We did not provide any monetary compensation for this annotation.
@@ Appendix I On-targetness
Appendix K Examples on Translating Sentences
While this work focuses on single word translation, we show examples in this sentence of applying iterative logit lens as described in § 3 for a sentence-level translation task. We observe that intermediate layers exhibit code-switching behavior that interpolates between the correct target language and English. We invite future work in studying the translation barrier for sequence-level tasks.
K.1 Example 1
@@
| Correct Translation | {CJK*}UTF8gbsn我的三分钟还没开始吧? |
| Layer 32 (Final Layer) | {CJK*}UTF8gbsn我的3分钟还没有开始,是吗? |
| Layer 29 | {CJK*}UTF8gbsn我的3 minute还没有开始,是吗? |
| Layer 27 | my three minutes chua chua{CJK*}UTF8gbsn开始吧 right{CJK*}UTF8gbsn吗? |
| Layer 25 | my three minutes already chua{CJK*}UTF8gbsn开始吧 right{CJK*}UTF8gbsn吧? |
K.2 Example 2
@@
| Correct Translation | {CJK*}UTF8gbsn我是说,我已经够紧张的了 |
| Layer 32 (Final Layer) | {CJK*}UTF8gbsn我的意思是,我已经很紧张了。 |
| Layer 29 | {CJK*}UTF8gbsn我的意思是我我 already足够紧张了。 |
| Layer 27 | {CJK*}UTF8gbsn我mean是我我 already cukup nervous了。 |
| Layer 25 | {CJK*}UTF8gbsn我 already là, already already already nervous了。 |
K.3 Example 3
@@
| Correct Translation | {CJK*}UTF8gbsn在加拿大,我们有很好的医保制度 |
| Layer 32 (Final Layer) | {CJK*}UTF8gbsn此外,在加拿大,我们有一个很棒的医疗系统。 |
| Layer 29 | {CJK*}UTF8gbsn此外,我们加拿大,我们拥有 fantast好的的健康 system。 |
| Layer 27 | {CJK*}UTF8gbsn此外,我们 Canada我们我们有 fantast好的的 healthcare system。 |
| Layer 25 | {CJK*}UTF8gbsn此外,我们 Canada we我们有 superb好的的 healthcare system。 |
K.4 Example 3

### v2

@@ Acknowledgments
We would like to thank Kaiser Sun and Krithika Ramesh for their helpful feedback. We would like to thank Jacob Wojnas, Varsini Sakthivadivel Ramasamy, and Suhas Sasetty, for providing annotations for the manual evaluation of the Polish, Tamil, and Telugu datasets respectively. This work is partially supported by the AI and Advanced Computing Institute at Schmidt Sciences.
@@ Appendix A Languages
| Language | Language (Sub)Family | FLORES+ code | Supported by Aya | Supported by Llama | Resource level |
| English | Germanic | eng_Latn | ✓ | ✓ | 7.0M |
| Cebuano | Austronesian | ceb_Latn | - | - | 6.1M |
| German | Germanic | deu_Latn | ✓ | ✓ | 3M |
| French | Romance | fra_Latn | ✓ | ✓ | 2.7M |
| Dutch | Germanic | nld_Latn | ✓ | - | 2.2M |
| Russian | East Slavic | rus_Cyrl | ✓ | - | 2.1M |
| Spanish | Romance | spa_Latn | ✓ | ✓ | 2M |
| Italian | Romance | ita_Latn | ✓ | ✓ | 1.9M |
| Polish | West Slavic | pol_Latn | ✓ | - | 1.7M |
| Chinese (Simplified) | Sinitic | zho_Hans | ✓ | - | 1.5M |
| Chinese (Traditional) | Sinitic | zho_Hant | ✓ | - | (see above) |
| Japanese | Japonic | jpn_Jpan | ✓ | - | 1.5M |
| Ukranian | East Slavic | ukr_Cyrl | ✓ | - | 1.4M |
| Vietnamese | Austroasiatic | vie_Latn | ✓ | - | 1.3M |
| Arabic | Semitic | arb_Arab | ✓ | - | 1.3M |
| Portuguese | Romance | por_Latn | ✓ | ✓ | 1.2M |
| Persian | Indo-Iranian | pes_Arab | ✓ | - | 1M |
| Catalan | Romance | cat_Latn | - | - | 800k |
| Indonesian | Austronesian | ind_Latn | ✓ | - | 700k |
| Korean | Koreanic | kor_Hang | ✓ | - | 700k |
| Turkish | Turkic | tur_Latn | ✓ | - | 600k |
| Czech | West Slavic | ces_Latn | ✓ | - | 600k |
| Romanian | Romance | ron_Latn | ✓ | - | 500k |
| Hebrew | Semitic | heb_Hebr | ✓ | - | 400k |
| Uzbek | Turkic | uzn_Latn | - | - | 300k |
| Greek | Hellenic | ell_Grek | ✓ | - | 300k |
| Tamil | Southern Dravidian | tam_Taml | - | - | 200k |
| Thai | Kra-Dai | tha_Thai | - | ✓ | 200k |
| Hindi | Indic | hin_Deva | ✓ | ✓ | 200k |
| Telugu | Southern Dravidian | tel_Telu | - | - | 100k |
| Swahili | Niger-Congo | swh_Latn | - | - | 100k |
| Marathi | Indic | mar_Deva | - | - | 100k |
| Bosnian | South Slavic | bos_Latn | - | - | 100k |
| Yoruba | Niger-Congo | yor_Latn | - | - | 30k |
| Nepali | Indo-Aryan | nep_Deva | - | - | 30k |
| Amharic | Semitic | amh_Ethi | - | - | 15k |
@@ Appendix B Dataset Details
In order to curate the source concepts that we translate to every language, we prompted ChatGPT to give us instances each of English nouns, verbs, adjectives and adverbs (totalling words). We then filtered out any words that did not appear in the English NLTK WordNet Bird and Loper (2004); Miller (1994), and manually vetted the remaining words to retain “general” concepts (i.e. excluding any proper nouns as well as highly specific cultural terms). We also checked for offensive words and did not encounter any. Finally, we sample words per part-of-speech of the filtered words, and conduct all our experiments on the resulting concepts.
We collect translations, including synonyms and alternate word senses, for these concepts for each of the other languages studied using Google Translate as mentioned in § 2. For average number of equivalent words per source word per language, see Table 3. We would like to note that the above method yielded only 1 equivalent per source word in some languages, like Chinese and Amharic and that this may have affected our results for these languages.
@@
We further asked proficient speakers of Hindi, Marathi, Uzbek, Telugu, and Tamil, to go through entries of our dataset, and report the number of entries with incorrect listed translations. Note that the last four languages are among the lowest-resourced in study. We report the resulting error rates in Table 4.
The consent of all annotators was obtained regarding the usage of the collected data for accuracy and coverage reporting of the lexica. We did not provide any monetary compensation for this annotation.
| Language | Error % |
| Hindi | 2% |
| Marathi | 7% |
| Uzbek | 0% |
| Tamil | 1% |
| Telugu | 2% |
@@ Appendix I On-targetness
Appendix K Arithmetic task
See examples of our input, intermediate responses, and the final generated responses in Table 6 and Table 7. We tried various prompting settings, including a zero-shot prompt. We found that both aya-23 and llama-3.1 tended to give sentence-long responses in those settings. The above format, i.e. 1-shot prompting with the example in the target language, worked best to constrain the answer.
We also tried a variant of this task where the input itself in provided in the target language. Our main observations were that even in this case, the intermediate layers exhibit answers in HRLs before switching to the target language. Further, this degrades performance for all languages, especially in Marathi and Swahili. This ties back to our discussion in § 4 about source language comprehensibility.
Appendix L Examples for sentence-level translation
While this work focuses on single word translation, we show examples of applying iterative logit lens as described in § 3 for a sentence-level translation task. We observe that intermediate layers exhibit code-switching behavior that interpolates between the correct target language and English. We invite future work in studying the translation barrier for sequence-level tasks.
L.1 Example 1
@@
| Correct Translation | 我的三分钟还没开始吧? |
| Layer 32 (Final Layer) | 我的3分钟还没有开始,是吗? |
| Layer 29 | 我的3 minute还没有开始,是吗? |
| Layer 27 | my three minutes chua chua开始吧 right吗? |
| Layer 25 | my three minutes already chua开始吧 right吧? |
L.2 Example 2
@@
| Correct Translation | 我是说,我已经够紧张的了 |
| Layer 32 (Final Layer) | 我的意思是,我已经很紧张了。 |
| Layer 29 | 我的意思是我我 already足够紧张了。 |
| Layer 27 | 我mean是我我 already cukup nervous了。 |
| Layer 25 | 我 already là, already already already nervous了。 |
L.3 Example 3
@@
| Correct Translation | 在加拿大,我们有很好的医保制度 |
| Layer 32 (Final Layer) | 此外,在加拿大,我们有一个很棒的医疗系统。 |
| Layer 29 | 此外,我们加拿大,我们拥有 fantast好的的健康 system。 |
| Layer 27 | 此外,我们 Canada我们我们有 fantast好的的 healthcare system。 |
| Layer 25 | 此外,我们 Canada we我们有 superb好的的 healthcare system。 |
L.4 Example 3
@@
Appendix M More Related Work
Cross-lingual Representations